OpenStudy (rayford):

x3over 4 divided by x2 over 3

OpenStudy (anonymous):

Is this the problem? \[\frac{x^3}{4} ÷ \frac{x^2}{3}\]

OpenStudy (rayford):

no the x is on the side

OpenStudy (rayford):

please im sorry but hurry my brother is trying to take the computer!

OpenStudy (anonymous):

Alright. Is it \[x\frac{3}{4} ÷ x\frac{2}{3}\]If it's not, can you draw it?

OpenStudy (rayford):

yes ur right

OpenStudy (anonymous):

Alright. I'm assuming these are mixed numbers here... \[\frac{4x + 3}{4} ÷ \frac{3x + 2}{3} \implies \frac{4x + 3}{4} \times \frac{3}{4x + 3}\]

OpenStudy (anonymous):

Oops. \[\frac{4x + 3}{4} \times \frac{3}{3x + 2}\]

OpenStudy (anonymous):

@apoorvk Can you take over here quickly? I have to go for a few minutes...

OpenStudy (rayford):

im sorry but thats incorrect it was x1/12

OpenStudy (apoorvk):

x1/12?? hey could you please use the drawing tool below to draw the question?